OpenStudy (anonymous):

a van accelerates at 3.20 m/s s^2 to the south. The net force acting on the van is 5.50 x 10^3 N What is the mass of the van?

OpenStudy (anonymous):

Force equals mass times acceleration aka F=ma. So, rearrange the equation to F/a=m, and then sub in the variables. So, in this case, it would be 5.5x 10^3 / 3.2 meters per second squared = 1718.75 kg.

OpenStudy (anonymous):

Assuming there is no friction, the only force acting in the horizontal direction will be 5.50 x 10^3 N. So Fnet = 5.50 x 10^3 N. Fnet = ma , where a = acceleration Therefore, m = Fnet/a Plug in the numbers and solve for m.